Calculus II, Spring 2009 (http://www.math.nthu.edu.tw/˜ wangwc/)

Homework Assignment for Week 14

Assigned May 26, 2009.

1. Section 15.1: Problems 27, 28.

2. Section 15.4: Problems 3, 6, 15, 17, 21, 22, 26, 27, 29.